8000 is invested in cds and stocks the cd pay annual interest rate of 3% and the stocks end up yielding an annual interest rate of 4% if the amount of interest received after one year is $275 then how much was invested into the cds
Answer by Boreal(15235)

x is cd
8000-x is stocks
.03x+.04(8000-x)=0.03x+320-.04x=275
-0.01x=-45
x=$4500 cds ANSWER and 3%=$135
8000-x=$3500 stocks @4%=$140
That total is $275